Assuming that Nations 1 and 2 are both large and starting from the equilibrium level of


national income and equilibrium in the trade balance in Nation 1 and given that 𝑀𝑃𝑆1 = 0.20; 𝑀𝑃𝑆2 = 0.15; 𝑀𝑃𝑀1 = 0.20 π‘Žπ‘›π‘‘ 𝑀𝑃𝑀2 = 0.10, find the change in the


equilibrium level of national income and the trade balance in Nation 1 for:


a) An autonomous increase in the exports of Nation 1 of 200 that replaces domestic


production in Nation 2


b) An autonomous increase in investment of 200 in Nation 1


c) An autonomous increase in investment of 200 in Nation 2.



1
Expert's answer
2022-02-28T11:34:10-0500

A)"K^{"} = \frac{1} { MPS1 + MPM1 + MPM2 (MPS1/MPS2)}"

= "\frac{1}{ 0.20 +0.20+ 0.10 (0.20/0.15)}"

="\frac{1}{0.533} = 1.88"

"\Delta{YE} =(\Delta{X})(K") = (200)\\times(1.88) = 376"

"\Delta{M} =(\Delta YE)(MPM_1)=(376)\times(0.20) = 75.2"

"\Delta S=(\Delta YE)(MPS_1)=376\times0.20"

"\Delta X = \Delta S + \Delta M = 75.2+ 75.2 = 150.4" so that

"\Delta X-\Delta M" = 75.2= Nation 1's trade surplus.

B)"K^{*} = \frac{1+ MPM2/MP52}{MPS1+ MPM1+ MPM2(MPS1/MPS)}"

= ("\frac{1+0.101}{0.15/0.533})"

="\frac{1.667}{0.533}" =3.13

"\Delta YE= (\Delta IX)= (200X3.13)" =626

"\Delta M=(\Delta YE\times MPM_1)=(626\times0.20)" =125.2

"\Delta S= (\Delta YE\times MPS1) =(626X0.20)" =125.2

200+"\Delta X" =125.2+125.2

and "\Delta X" = 50.4

so that"\Delta X-\Delta M" =50.4-125.2=74.8

c)"K^{*} = \frac{1+ MPM2/MP52}{MPS1+ MPM1+ MPM2(MPS1/MPS)}"

"(\frac{1+0.101}{0.20/0.533})"

"\frac{1.667}{0.106}=15.72"

"\Delta YE= (\Delta IX)= (200X15.72)=3144"

"\Delta M=(\Delta YEX\times MPM_1)=(626\times0.15)=93.9"

"\Delta S= (\Delta YE\times MPS1) =(626X0.15)=93.9"

200+"\Delta" X=93.9+93.9=187.8

"\Delta X=" 134.8

"\Delta X" -"\Delta M" =134.8-93.9=40.9